# Polarization Patterns of Non-Circular Hotspots around Kerr Black Holes: A Preliminary Study

### v2

@@ Abstract
The multi-wavelength polarized light signals from supermassive black holes have sparked many studies on polarized images of accretion disks and hotspots. However, the polarization patterns within the innermost stable circular orbit (ISCO) region remain to be explored. In this study, we focus on two specific types of orbits, namely the plunging geodesics inward from the ISCO and homoclinic geodesics, to uncover the polarization features associated with non-circular motion in a Kerr spacetime. For an on-axis observer, we specifically develop an approximate function to describe gravitational lensing along the azimuthal direction and establish a simplified synchrotron emission model. Based on these, we analyze the polarized patterns of hotspots accumulated over time and their Stokes parameters. Moreover, we explore the polarized image of the plunging region within a thin accretion disk.
1 Department of Physics, School of Physical Science and Technology, Ningbo University, Ningbo, Zhejiang 315211, China
@@ 1 Introduction
Optical observations of supermassive black holes have yielded significant results in recent years. Whether it be the millimeter-band images of Sagittarius A* (Sgr A*) and M87* captured by the Event Horizon Telescope (EHT) collaboration [1, 2] or the astrometry of near-infrared (NIR) flares from Sgr A* reported by the GRAVITY collaboration [3], both instances prompt a deeper exploration of black hole physics. The findings from EHT and GRAVITY encompass the information of polarization, offering more insights into the surrounding astronomical environments near the black holes [4, 5, 6]. The observed polarization angle, stemming from synchrotron radiation, encodes details about the plasma properties and magnetic field structures at the horizon-scale [7, 8, 9], aiding in the investigation of strong field physics [10, 11, 12, 13, 14].
@@
A simple but efficient way to analyze the NIR flares near Sgr A* is to consider these flares as time-varying images of a hotspot. This hotspot can be envisioned as a small, high-temperature, and neutral plasmoid, with its internal electrons emitting thermal synchrotron radiation [15, 16, 17, 18, 19, 20, 21]. One plausible mechanism for generating such a plasmoid is magnetic reconnection, a process in which magnetic field lines are severely curved and undergo reconnection to dissipate their energy [22, 23, 24, 25, 26]. The energy release in this process drives the plasma into high-energy plasmoids that are expelled from the reconnection region. The radiation characteristics of these plasmoids render them compelling candidates for the flares observed from Sgr A* [26].
@@
In theoretical research, it has been proposed a ring model, as outlined in [27] that the emissions originate from a circular ring encircling a Schwarzschild black hole. Viewing the ring as the accumulation of an orbiting hotspot over time, this model offers a simplified depiction of orbiting hotspots [28], where the locally defined magnetic field can be modulated to generate various optical appearances. Furthermore, by superimposing rings of different radii, the ring model can serve as a simplified representation of the accretion disk, facilitating the analysis of polarized images captured by the EHT [27]. The ring model was then expanded to encompass more general spherically symmetric spacetimes [29] and the rotating Kerr spacetime [30]. Using this analytic ring model, it has been shown in [31] that the twisty morphology in linear polarization is driven by the magnetic field structure in the emitting region, with the Doppler boosting and the lensing being subdominant factors. In [32], a more realistic accretion model near a Schwarzschild black hole was developed analytically. Additionally, investigations have delved into the impact of various types of spacetime on the polarization pattern within this ring model [33, 34, 35, 36].
@@
There exist several unexplored avenues in the current research on the polarization images. One of the most challenging tasks is to explore the horizon-scale magnetic field, which is intricately linked to the background magnetofluid [7]. Traditionally, the strength and the orientation of the magnetic field have been treated as phenomenological parameters in studies. Moreover, while the ring model adeptly captures the key characteristics of an orbiting hotspot or accretion disk, it is not well-suited for the domain within the Innermost Stable Circular Orbit (ISCO). The inner-ISCO region of an accretion flow, referred to as the plunging region, characterized by streamlines that can be roughly approximated by a cluster of plunging geodesics, could potentially make significant contributions to multi-wavelength observations. These observations could encompass the inner shadow observations by the EHT [37] and the X-ray spectra of stellar mass black holes [38, 39, 40]. For non-circular hotspots, particularly those perturbed from unstable orbits, the astrometry and the polarization patterns may diverge significantly from those of circular hotspots. This divergence could present new avenues for theoretical exploration into the flares observed from Sgr A* and potentially unlock novel insights into their nature.
@@
Recent studies have demonstrated that the trajectory functions of timelike geodesics perturbed from unstable circular orbits or the ISCO can be elegantly expressed using elementary functions [41, 42]. These findings have enabled the generation of trajectory images through purely analytical ways. In this work, we extend the methodology employed in studying circular orbit imaging and polarizations [27, 30] to encompass non-circular orbits within a Kerr spacetime. This extension utilizes the trajectory integrals outlined in [41, 42]. In our study, we position the distant observer along the axis of rotation of the black hole for simplicity, thereby deriving a set of analytical expressions for gravitational lensing. Our focus lies on the hotspots moving along two distinct types of trajectories: plunging geodesics and homoclinic orbits [41]. We analyze the polarized images accumulated over time for these hotspots, and carefully examine the variations in total intensity, the polarization vector along the trajectory, and the polarization Stokes parameters under diverse magnetic fields. We identify a range of polarization characteristics exhibited along different trajectories. Furthermore, we apply the aforementioned methodology to explore the polarized image of the plunging region within an accretion disk. We explore the distributions of total intensity and polarization vector across the disk plane.
@@ 2.2 Inspiraling orbits and homoclinic orbits
In this case, Eq. (2.10) can be integrated to yield the trajectory function parameterized by $r$ [44]
@@
where $f_{\pm}(r)$ is the same as in Eq. (2.2). These orbits are referred to as homoclinic orbits [41]. Physically, a homoclinic orbit corresponds to a hotspot forming at an unstable but bound circular orbit, and is subsequently perturbed away from the circular orbit. The right panel in Fig. 1 depicts a homoclinic orbit with parameters $a=0.2,r_{c}=r_{ib}+0.6M$, and $r_{3}\approx 11.96$. In this scenario, the hotspot is perturbed outwards from $r=r_{c}$, circles around it multiple times, and eventually reaches the “aphelion” at $r=r_{3}$. Subsequently, the hotspot retreats from the aphelion back towards $r=r_{c}$.
@@ 3 Imaging the hotspot
As our interest lies in the optical appearance of the hotspot, we need to compute the light propagations from the hotspot to a distant observer. In this section, we initially revisit the basic equations of light propagation in Kerr spacetime. Then, we discuss the map between the screen coordinates of the observer and the hotspot’s position, and provide a new fitting function of the map, which proves crucial for an analytical examination of the hotspot’s images. Subsequently, we introduce the setting for synchrotron radiation of the hotspot under external magnetic field.
@@ 3.2 Coordinates on the screen
The observations indicate small inclination angles for the supermassive black holes [1, 6, 8]. In this study, for simplicity, we restrict ourselves to an on-axis observer, $\theta_{o}=0^{\circ}$. From the angular potential, it is evident that only the photons with zero angular momentum, $\lambda=0$, can approach the spin axis. When a photon reaches the screen of the on-axis observer at the distance, its position of arrival is given by the polar coordinates [30]
@@
When deriving the above equations, the limit $r_{s}\gg M$ has been used. Despite this, Eq. (3.10) and Eq. (3.11) work well for nearly all values of $r_{s}>r_{+}$.
For an axisymmetric light source, such as a time-averaged accretion disk, Eq. (3.10) and Eq. (3.11) are sufficient to determine the optical appearance. This is because the emission profile does not depend on $\phi_{s}$, and the image does not depend on $\varphi$. However, when the hotspot serves as a point light source, a significant distortion along the $\phi$ direction occurs between the hotspot and its image on the screen, i.e., $|\varphi-\phi_{s}|$ could be large. This distortion originates from the frame-dragging effect on the light trajectory. Consequently, it is necessary to determine the map $\varphi(r_{s},\phi_{s})$ for hotspot imaging. For a photon emitted at $(r_{s},\phi_{s})$, the angle of arrival takes
@@
$\varphi-\phi_{s}\approx\int_{r_{s}}^{\infty}\frac{2Madr}{r^{2}\sqrt{r^{2}-r_{s}^{2}}}=\frac{2Ma}{r_{s}^{2}}\,.$ (3.13)
@@
$\varphi-\phi_{s}\approx\mathopen{}\mathclose{{}\left\{\begin{aligned} &\,\,\frac{2Ma}{r_{s}^{2}}\bigg{(}1-\frac{\tilde{\beta}}{(\frac{r_{s}}{r_{+}}-1)^{\tilde{\alpha}}}\ln{(\frac{r_{s}}{r_{+}}-1)}\bigg{)}\,,\quad\,\,r_{s}\textless r_{I}\,,\\ &\,\,\frac{2Ma}{r_{s}^{2}}\bigg{(}1+\frac{\tilde{\gamma}}{(\frac{r_{s}}{r_{+}}-1)}\bigg{)}\,,\quad\,\,r_{s}\geq r_{I}\,,\\ \end{aligned}}\right.$ (3.14)
@@
Fortunately, the approximation function Eq. (3.14) works well for a large parameter range of $a_{\star}$. In Fig. 2 we present the comparison between the approximation function Eq. (3.14) and the precise result of Eq. (3.12), under a small spin $a_{\star}=0.1$ and a high spin $a_{\star}=0.9$. The piecewise function is slightly discontinuous at the ISCO, where the largest error occurs. However, the error is rather small and can be neglected in subsequent discussions. By using Eq. (3.14), we can easily obtain the hotspot image for the on-axis observer, without the need to deal with lengthy elliptic integrals.
Figure 2: Comparsion between Eq. (3.12) and Eq. (3.14).
@@ 4.1 Results for inspiraling hotspots
Figure 3: From left to right, the columns present the trajectories, the polarization images, and the Q-U curve of the hotspot plunging from the prograde ISCO, under the vertical magnetic field Eq. (3.18). The constant $\mathcal{B}$ within the Papapetrou-Wald potential is set to be 1, without loss of generality. From top to bottom row, the black hole spin is set to be $a=0.1M,0.5M,0.9M$. For each case, the gray area denotes the region enclosed by $r_{+}$; the color bar represents the value of the total intensity; the polarization vector is represented by the red line segments on the screen; the Stokes parameters $Q,U$ are normalized by the maximal total intensity $I_{max}$ along the trajectory.
Figure 4: From left to right, the columns present the trajectories, the polarization images, and the Q-U curve of the hotspot plunging from the prograde ISCO, under the radial magnetic field Eq. (3.19), where the constant $\Psi$ is set to be 1, without loss of generality. From top to bottom, the rows show the black hole spin set to $a=0.1M,0.5M,$ and $0.9M$. For each case, the gray area denotes the region enclosed by $r_{+}$; the color bar represents the value of the total intensity; the polarization vector is represented by the red line segments on the screen; the Stokes parameters $Q,U$ are normalized by the maximal total intensity $I_{max}$ along the trajectory.
In Fig. 3 we depict the trajectory and the polarization image of the inspiraling hotspot under the vertical magnetic field Eq. (3.18), for various black hole spins. When depicting the images, we employ the Cartesian coordinates on the screen, $\alpha=\rho\cos{\varphi}\,,\beta=\rho\sin{\varphi}$. Due to the gravitational redshift, the total intensity exhibits a decreasing trend as the hotspot approaches the horizon. In the right colume of Fig. 3, the Stokes parameters along the hotspot’s trajectory is displayed on the Q-U plane, which we call the Q-U curve. Similar to the trajectory, the Q-U curve exhibits a right-handed spiral shape [^1] The chirality mentioned in this study is defined opposite to the direction of the observer’s line of sight.. From their definitions, ${\mathcal{Q}}_{o}+i\,{\mathcal{U}}_{o}={\mathcal{I}}_{o}\,e^{2i\chi_{net}+2i\varphi}$, it is evident that the phase variation of $Q,U$ is mainly determined by the hotspot’s motion in the $\phi$ direction.
@@
The detailed behaviors of the total intensity and the net EVPA are presented in Fig. 5. The region outside the ISCO is set to be filled with luminous hotspots moving along prograde circular geodesics, which have been well studied in the literature. It is easy to see the distinction in $I_{o}$ and $\chi_{net}$ outside and inside the ISCO. Note that in all four subplots, the connections at the ISCO are continuous. Under the vertical magnetic field, the net EVPA increases as the hotspot nears the horizon. For a higher black hole spin, $\chi_{net}$ displays a larger value at the prograde ISCO, but a smaller value at the horizon. However, the growth rate of $\chi_{net}$ is largely insensitive to the black hole spin, and in all cases the net EVPA exhibits a counterclockwise pattern, $\chi_{net}<\pi/2$, as shown in the top-right panel of Fig. 5.

### v3

# Polarization Patterns of the Hotspots Plunging into a Kerr Black Hole

@@ Abstract
The multi-wavelength polarized light signals from supermassive black holes have sparked many studies on polarized images of accretion disks and hotspots. However, the polarization patterns within the innermost stable circular orbit (ISCO) region remain to be explored. In this study, we focus on two specific types of orbits, namely the plunging geodesics inward from the ISCO and homoclinic geodesics, to uncover the polarization features associated with non-circular motion in a Kerr spacetime. For an on-axis observer, we specifically develop an approximate function to describe gravitational lensing along the azimuthal direction and establish a simplified synchrotron emission model. Based on these, we analyze the time-integrated polarized images of hotspots and their Stokes parameters. Moreover, we explore the polarized image of the plunging region within a thin accretion disk.
1 Institute of Fundamental Physics and Quantum Technology,
& School of Physical Science and Technology, Ningbo University, Ningbo, Zhejiang 315211, China
@@ 1 Introduction
Electromagnetic observations of supermassive black holes have yielded significant results in recent years. Whether it be the millimeter-band images of Sagittarius A* (Sgr A*) and M87* captured by the Event Horizon Telescope (EHT) [1, 2] or the astrometry of near-infrared (NIR) flares from Sgr A* reported by the GRAVITY collaboration [3], the observations prompt a deeper exploration of black hole physics. The findings from EHT and GRAVITY encompass the information of polarization, offering more insights into the surrounding astronomical environments near the black holes [4, 5, 6]. The observed polarization angles, stemming from synchrotron radiations, encode the details about the plasma properties and magnetic field structures at the horizon-scale [7, 8, 9], aiding in the investigation of strong field physics [10, 11, 12, 13, 14].
@@
The study of NIR flares near Sgr A* is of particular importance in understanding the dynamics of the Galactic Center. A simple yet effective approach to analyze these flares is to model them as time-varying images of a hotspot. This hotspot is a compact, high-temperature, neutral plasmoid, whose internal electrons emit synchrotron radiations [15, 16, 17, 18, 19, 20, 21]. One plausible mechanism for generating such a plasmoid is via magnetic reconnection, a process in which magnetic field lines are severely curved and undergo reconnections to dissipate their energy [22, 23, 24, 25, 26]. The released energy drives the plasmoids and expels them from the reconnection region. The radiation characteristics of these plasmoids render them the compelling candidates for the flares [26]. The observations of polarized light curves at both NIR and millimeter wavelengths support the hotspot model for Sgr A* [16, 19, 27, 28, 6].
@@
In analytical research, a ring model has been proposed to study the polarized images of hotspots around a Schwarzschild black hole, as outlined in [29]. This model assumes an emitting circular ring, interpreted as the accumulation of an orbiting hotspot over time [30]. By modulating the locally defined magnetic field, one can generate a variety of polarizations in this model. By superimposing rings of different radii, the model can also produce a simplified version of the polarized image of an accretion disk. The ring model has since been extended to more general spherically symmetric spacetimes [31] and the Kerr spacetime [32]. Using this analytic model, it has been demonstrated in [33] that the twisty morphology observed in linear polarization is primarily dictated by the magnetic field structure, with the Doppler boosting and the gravitational lensing playing subdominant roles. Additionally, a more realistic analytical model of accretion near a Schwarzschild black hole was developed in [34]. Further investigations have explored the influence of different spacetime geometries on the polarization patterns predicted by the ring model [35, 36, 37, 38].
@@
There exist several unexplored avenues in the current research on the polarized images. One of the most challenging tasks is to explore the horizon-scale magnetic field, which is intricately linked to the background magnetofluid [7]. Traditionally, the strength and the orientation of the magnetic field have been treated as phenomenological parameters in studies. Moreover, while the ring model captures the basic characteristics of orbiting hotspots or disk emission, it is not well-suited for the region within the Innermost Stable Circular Orbit (ISCO). The inner-ISCO region of an accretion flow, referred to as the plunging region, characterized by the streamlines that can be roughly approximated by a cluster of plunging geodesics, could make significant contributions to multi-wavelength observations. These observations could encompass the inner shadow observations by the EHT [39] and the X-ray spectra of stellar mass black holes [40, 41, 42]. Furthermore, for non-circular hotspots, particularly those perturbed from unstable orbits, their astrometry and polarization patterns may diverge largely from those of circular hotspots. This divergence presents new avenues for theoretical exploration into the flares observed from Sgr A* and potentially unlock novel insights into their nature.
@@
Previous studies have shown that the trajectories of timelike geodesics perturbed from unstable circular orbits or the ISCO can be expressed in terms of elementary functions [43], enabling the purely analytical generation of hotspot images. In this work, we extend the study of imaging and polarization of circular orbits [29, 32] to non-circular orbits in a Kerr spacetime. This extension utilizes the trajectory integrals outlined in [43, 44]. For simplicity, we place the distant observer along the black hole’s rotation axis in most cases, allowing us to derive analytical expressions for gravitational lensing. Our focus is on hotspots moving along two distinct types of trajectories: plunging geodesics and homoclinic orbits [43]. We analyze the time-averaged polarized images of these hotspots, carefully examining variations in total intensity, the polarization vector along the trajectory, and the Stokes parameters under different magnetic field configurations. We identify a range of polarization characteristics associated with different trajectories. Furthermore, we apply this methodology to study the polarized image of the plunging region within an accretion disk, exploring the distributions of total intensity and polarization vectors across the disk plane.
@@ 2.2 Inspiraling orbits and homoclinic orbits
In this case, Eq. (2.10) can be integrated to yield the trajectory function parameterized by $r$ [46, 47]
@@
where $f_{\pm}(r)$ is the same as in Eq. (2.2). These orbits are referred to as homoclinic orbits [43, 48]. Physically, a homoclinic orbit corresponds to a hotspot forming at an unstable but bound circular orbit, and is subsequently perturbed away from the circular orbit. The right panel in Fig. 1 depicts a homoclinic orbit with parameters $a=0.2,r_{c}=r_{ib}+0.6M$, and $r_{3}\approx 11.96$. In this scenario, the hotspot is perturbed outwards from $r=r_{c}$, circles around it multiple times, and eventually reaches the “aphelion” at $r=r_{3}$. Subsequently, the hotspot retreats from the aphelion back towards $r=r_{c}$.
@@ 3 Imaging the hotspot
As our interest lies in the images of the hotspot, we need to compute the light propagations from the hotspot to a distant observer. In this section, we initially revisit the basic equations of light propagation in Kerr spacetime. Then, we discuss the map between the screen coordinates of the observer and the hotspot’s position, and provide a new fitting function of the map, which proves crucial for an analytical examination of the hotspot’s images. Subsequently, we introduce the setting for synchrotron radiation of the hotspot under external magnetic field.
@@ 3.2 Coordinates on the screen
The observations indicate small inclination angles for the supermassive black holes [1, 6, 8]. In this study, for simplicity, we restrict ourselves to an on-axis observer. From the angular potential, it is evident that only the photons with zero angular momentum, $\lambda=0$, can approach the spin axis. When a photon reaches the screen of the on-axis observer at the distance, its position of arrival is given by the polar coordinates [32]
@@
When deriving the above equations, the limit $r_{s}\gg M$ has been used. However, the above equations work quite well even at the horizon scale ($r_{s}\sim r_{+}$) for all Kerr spin, as demonstrated in the Fig.5 in [51]. Therefore, we can safely use Eq. (3.10) and Eq. (3.11) to study the images of a hotspot close to the event horizon.
Figure 2: Comparision between Eq. (3.12) and Eq. (3.14) for two different Kerr parameters $a=0.1M$ and $a=0.9M$. In the upper plots, the solid lines represent the exact results, $\Delta\varphi_{\text{int}}$, obtained by numerically integrating Eq.(3.12), while the dashed lines correspond to the approximate results, $\Delta\varphi_{\text{fit}}$, computed using the fitted expression in Eq.(3.14). The lower plot illustrates the deviations, defined as $\delta_{\Delta\varphi}\equiv\Delta\varphi_{\text{int}}-\Delta\varphi_{\text{fit}}$.
For an axisymmetric light source, such as a time-averaged accretion disk, Eq. (3.10) and Eq. (3.11) are sufficient to determine the image. This is because the emission profile does not depend on $\phi_{s}$, and the image does not depend on $\varphi$. However, when the hotspot serves as a point light source, a significant distortion along the $\phi$ direction occurs between the hotspot and its image on the screen, i.e., $|\varphi-\phi_{s}|$ could be large. This distortion originates from the frame-dragging effect on the light trajectory. Consequently, it is necessary to determine the map $\varphi(r_{s},\phi_{s})$ for hotspot imaging. For a photon emitted at $(r_{s},\phi_{s})$, the angle of arrival takes
@@
$\Delta\varphi\equiv\varphi-\phi_{s}\approx\int_{r_{s}}^{\infty}\frac{2Madr}{r^{2}\sqrt{r^{2}-r_{s}^{2}}}=\frac{2Ma}{r_{s}^{2}}\,.$ (3.13)
@@
$\Delta\varphi\approx\mathopen{}\mathclose{{}\left\{\begin{aligned} &\,\,\frac{2Ma}{r_{s}^{2}}\mathopen{}\mathclose{{}\left[1-\frac{\tilde{\beta}}{(\frac{r_{s}}{r_{+}}-1)^{\tilde{\alpha}}}\ln{(\frac{r_{s}}{r_{+}}-1)}}\right]\,,\quad\,\,r_{s}\textless r_{I}\,,\\ &\,\,\frac{2Ma}{r_{s}^{2}}\mathopen{}\mathclose{{}\left[1+\frac{\tilde{\gamma}}{(\frac{r_{s}}{r_{+}}-1)}}\right]\,,\quad\,\,r_{s}\geq r_{I}\,,\\ \end{aligned}}\right.$ (3.14)
@@
Fortunately, the approximation function Eq. (3.14) works well for a large parameter range of $a_{\star}$. In Fig. 2 we present the comparison between the approximation function Eq. (3.14) and the precise result of Eq. (3.12), for a small spin $a_{\star}=0.1$ and a high spin $a_{\star}=0.9$. We also plot the deviation from the precise result, $\delta_{\Delta\varphi}=\Delta\varphi_{\text{int}}-\Delta\varphi_{\text{fit}}$. The deviation function $\delta_{\Delta\varphi}$ is slightly discontinuous at the ISCO, where the largest error occurs. For $a=0.9M$, the maximum error is $\delta_{\Delta\varphi\text{max}}<0.03\approx 1.72^{\circ}$, while for $a=0.1M$, the maximum error is even smaller. Therefore, the error in using the approximation function is rather small and can be neglected in subsequent discussions. By using Eq. (3.14), we can easily obtain the hotspot images for the on-axis observer, without the need to deal with lengthy elliptic integrals.
@@ 4.1 Results for inspiraling hotspots
Figure 3: From left to right, the columns present the trajectories, time-integrated polarized images, and the Q-U curve of the hotspot plunging from the prograde ISCO, under the vertical magnetic field Eq. (3.18). The constant $\mathcal{B}$ within the Papapetrou-Wald potential is set to be 1, without loss of generality. From top to bottom row, the black hole spin is set to be $a=0.1M,0.5M,0.9M$. For each case, the gray area denotes the projection of the black hole; the color bar represents the value of the total intensity; the polarization vector is represented by the red line segments on the screen; the Stokes parameters $Q,U$ are normalized by the maximal total intensity $I_{max}$ along the trajectory.
Figure 4: From left to right, the columns present the trajectories, time-integrated polarized images, and the Q-U curve of the hotspot plunging from the prograde ISCO, under the radial magnetic field Eq. (3.19), where the constant $\Psi$ is set to be 1, without loss of generality. From top to bottom, the rows show the black hole spin set to $a=0.1M,0.5M,$ and $0.9M$. For each case, the gray area denotes the projection of the black hole; the color bar represents the value of the total intensity; the polarization vector is represented by the red line segments on the screen; the Stokes parameters $Q,U$ are normalized by the maximal total intensity $I_{max}$ along the trajectory.
In Fig. 3 we depict the trajectory and the time-integrated polarized image of the inspiraling hotspot under the vertical magnetic field Eq. (3.18), for various black hole spins. When depicting the images, we employ the Cartesian coordinates on the screen, $\alpha=\rho\cos{\varphi}\,,\beta=\rho\sin{\varphi}$. Due to the gravitational redshift, the total intensity exhibits a decreasing trend as the hotspot approaches the horizon. In the right colume of Fig. 3, the Stokes parameters along the hotspot’s trajectory is displayed on the Q-U plane, which we call the Q-U curve. Similar to the trajectory, the Q-U curve exhibits a right-handed spiral shape [^1] The chirality mentioned in this study is defined opposite to the direction of the observer’s line of sight.. From their definitions, ${\mathcal{Q}}_{o}+i\,{\mathcal{U}}_{o}={\mathcal{I}}_{o}\,e^{2i\chi_{net}+2i\varphi}$, it is evident that the phase variation of $Q,U$ is mainly determined by the hotspot’s motion in the $\phi$ direction.
@@
The detailed behaviors of the total intensity and the net EVPA are presented in Fig. 5. The region outside the ISCO is set to be filled with luminous hotspots moving along prograde circular geodesics, which have been well studied in the literature. It is easy to see the distinction in $I_{o}$ and $\chi_{net}$ outside and inside the ISCO. Note that in all four subplots, the connections at the ISCO are continuous. Under the vertical magnetic field, the net EVPA increases as the hotspot nears the horizon. For a higher spin, $\chi_{net}$ displays a larger value at the prograde ISCO, but a smaller value at the horizon. However, the growth rate of $\chi_{net}$ is largely insensitive to the spin, and in all cases the net EVPA exhibits a counterclockwise pattern, $\chi_{net}<\pi/2$, as shown in the top-right panel of Fig. 5.